The final exam scores of students taking a biology course at a university are normally distributed with a population mean of 73 and population standard deviation of 15. Nine students are randomly selected from the university and let x be their average score.
What is the probability that their average biology score x is between 60 and 80?

Here's what I have so far...
z=(60-73)/5, z=(80-73)/5
z=-2.6;1.4
P(-2.6 < x < 1.4)
P(z<1.4)-P(z<-2.6)

Answer by stanbon(75887)

The final exam scores of students taking a biology course at a university are normally distributed with a population mean of 73 and population standard deviation of 15. Nine students are randomly selected from the university and let x be their average score.
What is the probability that their average biology score x is between 60 and 80?
Here's what I have so far...
z=(60-73)/5,e z=(80-73)/5
z=-2.6;1.4
P(-2.6 < x < 1.4)
P(z<1.4)-P(z<-2.6)
----------------------
You say the standard deviation is 15,
but you use 5 as the standard deviation.
----
Assuming 5 is the correct number, you
z-values are correct.
----
Your conclusion should be the following:
P(60 < x < 80) = P(-2.6 < z < 1.4) = 0.9146
==============================================
Cheers,
Stan H.
=============